Please help or confirm my answer.
If f(x)= x^2-x+7, find f(x-2)
My answer is (x-2)^2 =4, -x(x-2)= +2, then just +7
4x + 2 +7
x=9
Thanks

Answer by Alan3354(69443)

Please help or confirm my answer.
If f(x)= x^2-x+7, find f(x-2)
My answer is (x-2)^2 =4, -x(x-2)= +2, then just +7
4x + 2 +7
x=9
----------------------
If f(x)= x^2-x+7, find f(x-2)
Sub x-2 for x
f(x-2) = x^2-4x+4 -x+2 + 7
= x^2 - 5x + 11
That's all
-------------
You don't solve it, just substitute.